OpenStudy (anonymous):

The Mecklanberg public school district stipulates that a standard school day will average at least 5.5 hr excluding meal breaks. For the first 4 days of one school week, bad weather resulted in school days of 4 hr, 6.5 hr, 3.5 hr, 6.5 hr. How long must the friday school day be in order to average at least 5.5 hr for the week? 4, 6.5, 3.5, 6.5, + x = 20.5+ x /5=5.5 =20.5+ x = 27.5 x = 7 or 5 working days, total working hours 5.5 =5.5x5=27.5 =4+6.5+3.5+6.5+friday=20.5 + friday =27.5-20.5=7 hrs x=7 hrs. Is this problem going to be an equation or an inequality? And how would I

OpenStudy (anonymous):

How would I write it?

OpenStudy (anonymous):

1st aproach is coreect

OpenStudy (anonymous):

would it be an equation or equality

OpenStudy (anonymous):

it is eqation

OpenStudy (kropot72):

Let Friday hours = F total required hours for week = 5 * 5.5 = 4 + 6.5 + 3.5 + 6.5 + F F = 27.5 - 20.5 = 7 hours

OpenStudy (anonymous):

so kropot72, would it be an equation

OpenStudy (kropot72):

The answer is just a number followed by the units (hours). A series of equations is needed to fully show how the answer was obtained.

OpenStudy (anonymous):

Great. Thank you.

OpenStudy (kropot72):

You're welcome :)